OpenStudy (highschoolmom2010):

To landscape her 70 ft-by-60 ft rectangular backyard, your aunt is planning first to put down a 4-in. layer of topsoil. She can buy bags of topsoil at $2.50 per 3-ft^3 bag, with free delivery. Or, she can buy bulk topsoil for $22.00/yd^3, plus a $20 delivery fee. Which option is less expensive? Explain.

OpenStudy (anonymous):

total area= 70*60 1st option=> 70*60/3 2nd option=> 70*60/22 +20 convert yard firstly

OpenStudy (phi):

My first thought is figure out how many cubic feet of soil you need. that means you should change the depth of 4 inches to 1/3 of a foot, and figure the volume of a rectangular prism.

OpenStudy (highschoolmom2010):

V=bh V=(70*60)*1/3 V=4200*1/3 V=1400 ft^3 i think

OpenStudy (phi):

yes. Now figure out how much option A costs.

OpenStudy (highschoolmom2010):

1400 ft^3*$2.50=3500$

OpenStudy (phi):

I think the bags contain 3 cu ft of top soil

OpenStudy (highschoolmom2010):

1400/3=467 467*2.50=1167.5

OpenStudy (phi):

yes. now option B. I would change cubic yards into cubic feet so we can get a dollars per cubic foot price from $22 per cubic yard. a yard is 3 feet. a cubic yard is a 3 foot by 3 foot by 3 foot cube. You need to find the volume of a cube with a side of 3 feet.

OpenStudy (highschoolmom2010):

v=bh v=9*3 V=27

OpenStudy (phi):

yes there are 27 ft^3 in 1 yd^3 so now you have $22 per 27 ft^3 how much does it cost for 1400 ft^3 ? and then add in the $20 delivery charge.

OpenStudy (phi):

I would do \[ \frac{$22}{27 \text{ ft}^3 } \cdot 1400 \text{ ft}^3 +$20\]

OpenStudy (highschoolmom2010):

1160.74

OpenStudy (phi):

Now the final part. Which option is less expensive?

OpenStudy (highschoolmom2010):

2nd one